OpenStudy (daisiablis11):

How is tan^-1(1) = pi/4?

satellite73 (satellite73):

for the same reason that \[\tan(\frac{\pi}{4})=1\]

satellite73 (satellite73):

arctangent is the number (angle if you like) between \(-\frac{\pi}{2}\) and \(\frac{\pi}{2}\) whose tangent is given you are given that the tangent is 1, so the angle must be \(\frac{\pi}{4}\)

OpenStudy (daisiablis11):

Thanks again. I should really review my unit circle and algebra.

satellite73 (satellite73):

yw (good idea)

OpenStudy (daisiablis11):

I thought open study was closing btw?